# SiM3D: Single-instance Multiview Multimodal and Multisetup 3D Anomaly Detection Benchmark

### v1

@@ 2 Related Work
ADS Methods. Following the setup defined in the MVTec-AD benchmark, most ADS methods [22] are designed to estimate 2D anomaly maps from a single image [3, 43, 19, 29, 26, 13, 37, 27, 27, 39, 44, 34, 40, 20, 38, 23, 41, 31, 28, 16, 9, 15, 10, 2]. One of the most popular methods is PatchCore [30], which uses feature extractors pre-trained on large datasets [8, 24, 17] to gather features from nominal samples into a memory bank. At test time, input sample features are compared to this bank to detect anomalies. Another more recent yet popular approach, EfficientAD [1], uses a lightweight feature extractor and a student-teacher model to identify anomalies based on feature prediction discrepancy. The introduction of the MVTec 3D-AD benchmark has led to the development of several multimodal approaches that use images and 3D data to generate 2D anomaly maps [18, 36, 14, 32]. Among them, BTF [18] is a memory bank approach inspired by PatchCore, in which each element is the concatenation of 2D and 3D features extracted by a pre-trained image backbone and a traditional point cloud descriptor, i.e., FPFH [33], respectively. M3DM [36] improves BTF using powerful transformer-based pre-trained backbones, such as DINO-v1[8] and Point-MAE [25], to extract features from RGB images and point clouds. It also learns to fuse 2D and 3D features into multimodal features. During inference, a trained One-Class SVM learns to aggregate modality-specific scores. Recently, starting from the same transformer-based backbones as in [36], CFM [14] learns to map features from one modality to the other by nominal samples and then, at inference time, detects anomalies by pinpointing inconsistencies between observed and predicted features. In this paper, we employ the above-mentioned methods to investigate the challenges of SiM3D by adapting them to work in a multiview setup.
@@ 3.1 Material preparation and acquisition setup
Sensor. We acquired our dataset with a top-tier 3D scanner, the Atos Q from ZEISS (Fig. 3 (2.i)). It consists of a stereo pair with two 12 Mpx greyscale cameras and a light projector. We consider the left camera to be our reference. The images have a resolution of $4096\times 3000$ pixels. The point clouds can feature up to 12 million elements, with a point-to-point distance between 0.04 and 0.15 mm. The sensor is mounted on an industrial robot arm with high repeatability and precision (Fig. 3 (2.iii)).
Calibration. We calibrate the greyscale cameras with a standard chequerboard pattern, estimating the intrinsic parameters matrix, $A$, and the radial and tangential distortion coefficients, $\delta=[k_{1},k_{2},k_{3},p_{1},p_{2}]$. We also calibrate the rotation, $R_{pc}$, and translation, $T_{pc}$, from the point cloud to the left camera reference frame, with a precise dotted calibration pattern built by ZEISS (Fig. 3 (2.ii)) that provides centre coordinates with precision $<1$ mm. More details are in the supplementary.
3.2 Data Gathering
@@
3.3 Labelling
Since some anomalies, such as paintings and minor scratches, are visible only in greyscale images, while others, such as dents, are better visible in 3D, we must consider both sensing modalities to obtain precise 3D segmentation ground-truths for anomalous samples. Therefore, we develop a two-step labelling strategy, illustrated in Fig. 3 (4), that operates on both 2D images and 3D data. First, we manually annotate all the view images, $I_{i}$, where anomalies are clearly visible, obtaining dense 2D segmentation masks. We employ a different ID and colour for every defect to ease the voxel AUPRO computation (see Sec. 4). Subsequently, for each annotated image, $I_{i}$, we leverage the associated integrated mesh, $M$, the intrinsic parameters $A$, the point cloud to camera reference frame transformation, $(R_{\text{pc}},T_{\text{pc}})$, and the transformation between $v_{i}$ and $v_{1}$, $(R_{i},T_{i})$, to project the 2D annotations onto the integrated mesh. Then, since the projection may not be precise due to occlusions and slanted surfaces, each ground-truth mesh is thoroughly inspected and manually refined using a 3D visualisation and editing software, i.e., CloudCompare [12]. The labelling process was conducted by a team of 4 experts, with a final cross-check of the results. Finally, the annotated mesh is converted into a voxel grid with a voxel size of 2 mm. The choice of the size is determined by the smallest defect size discernible in the meshes. In the conversion process, a voxel is labelled as anomalous if it intersects at least one triangle labelled as such in the annotated mesh.
@@ 3.3 Labelling
Plastic Stool $35\times 35\times 30$ $330\times 326\times 317$ 22 1 real or 1 synth 10 10 12
Rubbish Bin $26\times 21\times 33$ $333\times 339\times 326$ 42 1 real or 1 synth 20 20 12
Wicker Vase $17\times 17\times 15$ $329\times 309\times 290$ 22 1 real or 1 synth 10 10 12
Bathroom Furniture $33\times 33\times 50$ $371\times 392\times 363$ 20 1 real or 1 synth 8 10 36
Container $20\times 25\times 10$ $327\times 308\times 286$ 94 1 real or 1 synth 46 46 12
Plastic Vase $12\times 12\times 9$ $326\times 308\times 286$ 99 1 real or 1 synth 48 49 12
Wooden Stool $48\times 42\times 45$ $331\times 367\times 327$ 15 1 real or 1 synth 6 7 12
Sink Cabinet $44\times 25\times 50$ $342\times 349\times 357$ 19 1 real or 1 synth 9 8 36
@@ 4.1 Dataset, Tasks, and Metrics
Figure 5: Qualitatives. Selected views with defects (left) and Anomaly Volumes for several methods (right) downsampled for visualization.
@@
Metrics. Akin to standard practice in 2D ADS [4], we evaluate detection performance by the Area Under the Receiver Operator Curve (AUROC) of the global anomaly scores predicted by the methods. As we compute a global score for each object instance, we name it instance-level AUROC (I-AUROC). Regarding segmentation performance, we evaluate methods with an adaptation of the popular Area Under the Per-Region Overlap (AUPRO) curve. First of all, we define the voxel-based Per-Region Overlap (PRO) metric as:
$\text{PRO}_{t}=\frac{1}{N}\sum_{n=1}^{N}\frac{|\overline{\Psi}_{t}\cap V_{n}|}{|V_{n}|}$ (1)
where $\overline{\Psi}_{t}$ is an Anomaly Volume binarised using a threshold, $t$, $N$ is the number of blobs, i.e., defects, in the ground-truth, and $V_{n}$ are the voxels belonging to that blob. Since during labelling, we annotated each defect in our dataset with a different ID (see Sec. 3.3), the $N$ defect blobs in a ground-truth are easily identifiable. This metric is computed across multiple thresholds, $t$, estimated at different False Positive Rates (FPRs). We calculate the FPR using only non-empty voxels. We sample FPRs from 0 to the maximum value uniformly, we estimate the corresponding threshold $t$, and we binarise the score grid accordingly. In this way, we construct a curve that plots the PRO values against the corresponding FPRs. The segmentation performance is obtained by integrating such a curve up to a specific False Positive Rate. As proposed in [14], we use 1% as integration bound for the PRO curve to set a challenging bar representative of the requirements of real industrial applications. Finally, we normalise the resulting area to the interval $[0,1]$. We dub this metric voxel-level AUPRO (V-AUPRO).
@@ 4.2 Baselines
Extending singleview methods to Multiview 3D ADS. Current ADS methods predict a 2D anomaly map from a singleview input, either unimodal, e.g., images, or multimodal, e.g., images and point clouds. To obtain a set of baselines for our novel multiview 3D ADS task, we extend popular and effective singleview methods by means of the 3D aggregation strategy illustrated in Fig. 4. Given the considered method, we process the data corresponding to an input view, $v_{i}$, so as to obtain a 2D Anomaly Map. Similarly to Sec. 3.3, given the mesh, $M$, intrinsic parameters $A$, point cloud to camera reference frame transformation, $(R_{\text{pc}},T_{\text{pc}})$ and view pose $v_{0}$, $(R_{i},T_{i})$, we can project each pixel of the 2D Anomaly Map to its corresponding 3D position. We discretise the 3D position, finding the corresponding voxel index. We associate that voxel with the corresponding anomaly score. We repeat the process for each view, keeping track of the scores that project onto the same voxels. Eventually, we compute the maximum among all scores projected into each voxel, obtaining the final Anomaly Volume. To extract the global score required by the I-AUROC, we compute the maximum across all the scores within the Anomaly Volume.
@@
Selected Methods. We extend some of the most popular and effective unimodal and multimodal singleview ADS methods, starting from their official code, and evaluate them on SiM3D. Regarding unimodal methods, we select PatchCore [30], with either WideResNet101 [42] or DINO-v2 [24] as backbone, to extract features from images of all views, $\{I_{i}\}_{i=1}^{N}$ of the single-instance training set to create memory banks. Moreover, we use EfficientAD [1] using all views as their training set. As for multimodal ADS methods, we use BTF [18] employing WideResNet101 [42] as the image encoder and FPFH [33] as the algorithm to extract point cloud features. We create memory banks using features extracted from images, $\{I_{i}\}_{i=1}^{N}$, and point clouds, $\{P_{i}\}_{i=1}^{N}$, of all views of the single-instance training set. We use the same strategy with M3DM [36]. As 2D feature extractor, we exploit DINO-v2. As point cloud feature extractor, instead of using a transformer backbone, i.e., Point-MAE [17], we utilise FPFH to extract 3D features, similarly to BTF. Indeed, we would need a severe downsample (from $\sim$5 – 7M to $\sim$8k points) to work with Point-MAE, and we verified that at such low resolution, most of the defects no longer show up in the point clouds. Finally, we select CFM [14], using all views and clouds as a single training set. Based on the same consideration as for M3DM, we use DINO-v2 and FPFH to produce 2D and 3D features. Due to computational constraints, all methods process images padded and downsampled to $1540\times 1540$. Moreover, for compatibility with the employed 2D feature extractors, which accept RGB images, we create 3-channel inputs by stacking our greyscale images thrice.
@@
3D Data pre-processing. To obtain less noisy data and improve the results of the above-mentioned baselines, we remove the background from 3D data. We estimate a 3D plane using a RANSAC-based plane segmentation algorithm, which, iteratively, randomly selects a small subset of points (e.g., 10) to fit a candidate plane model. For the candidate planes, we compute the consensus set with threshold $\tau$ on the point-to-plane distance. After 1000 iterations, we return the best-fitting plane. To account for spurious background artefacts, we shift the plane by an offset $\alpha$ and finally remove all points below the shifted plane. Given the different sizes of the objects, for each object type, we select different optimal thresholds $\tau$ and offsets $\alpha$, as detailed in the supplementary material. We apply this background removal technique to the integrated mesh, $M$, due to the larger number of background points, which facilitates the RANSAC algorithm. Then, from the background-filtered integrated mesh, we leverage ray-tracing techniques and known camera parameters to render, for each view, depth maps, $\{D_{i}\}_{i=1}^{N}$, pixel-aligned to the corresponding images, $\{I_{i}\}_{i=1}^{N}$ at the sensor resolution ($4096\times 3000$). Finally, these depth maps are padded and downsampled to $1540\times 1540$, and the depth is lifted in 3D via the camera parameters to obtain dense and clean point clouds. The depth maps and point clouds obtained by rendering the pre-processed meshes serve as the input data for all the experiments presented in the subsequent section and will be available in the SiM3D dataset.
@@ 5 Benchmark Results
Main Results. In  Tab. 3 and  Tab. 4 we report the results obtained on the SiM3D benchmark by extending the considered methods as Multiview 3D ADS baselines. By looking at the mean detection (Tab. 3) and segmentation (Tab. 4) performance, we notice how methods that natively process only RGB images outperform natively multimodal ones in both the real2real and synth2real setups. We posit that this is due to the main multimodal methods having been designed and tuned to process point clouds of the resolution featured by existing multimodal benchmarks, which is much lower than that of SiM3D, i.e. $\sim$ 8K vs $\sim$ 2.3M points (after downsampling). Hence, our experiments suggest that current multimodal ADS methods do not scale to significantly higher resolutions. In turn, we could not even deploy the native 3D feature extractors of M3DM and CFM, the methods currently excelling on [6, 7]. Furthermore, we notice how methods relying on memory banks, such as PatchCore and M3DM, tend to attain slightly better mean detection and segmentation performance than their counterparts within the same modality that require substantial training, i.e., EfficientAD and CFM. We argue that this may be ascribed to the single-instance setup, which renders the latter category of approaches harder to optimise due to the limited number of training images, which are as many as the views. The method most effectively extended to our multiview 3D task is Patchcore, which, equipped with either WRN-101 or DINO-v2, can achieve the best detection and segmentation performance in both setups. Indeed, it yields an I-AUROCs of 0.754 (WRN-101) and 0.540 (DINO-v2), a V-AUPRO@1% of 0.671 (DINO-v2) and 0.600 (WRN-101), in the real2real and synth2real setups, respectively. It is also worth pointing out how, due to the domain shift, we can notice a large gap between the real2real and synth2real setups in both best detection (0.754 vs. 0.540, I-AUROC) and segmentation performance (0.671 vs. 0.600, V-AUPRO@1%). The qualitative results, shown in Fig. 5, visually support the above observations. Overall, the relatively low, definitely far from saturated, performance figures provided by baselines designed by straightforward extension of state-of-the-art methods delivering impressive results in singleview 2D benchmarks hint at the novel and challenging nature of the task set forth by SiM3D. Thus, our experimental findings by baseline approaches reveal the need for more advanced and specific methods and paradigms, both in the real2real and, even more so, the synth2real setups.
@@
Alternative 3D Representations. Given the lack of feature extractors amenable to processing high-resolution point clouds, we explore the use of depth maps as the input modality to capture 3D cues. The 2D grid structure of depth maps allows for deploying popular backbones pre-trained on RGB data that support high-resolution inputs, such as DINO-v2. Thus, we create and evaluate additional baselines using depth maps and report the results in  Tab. 5. Similarly to greyscale images, we tile depth maps along the channel axis to ensure compatibility with the employed backbones. The results show that the use of depth maps rather than point clouds in multimodal methods tends to improve detection and segmentation performance by sizable margins, e.g., in the real2real setup, BTF yields 0.707 I-AUROC and 0.609 V-AUPRO when processing depths compared to 0.478 I-AUROC and 0.444 V-AUPRO when fed with point clouds (row 5 vs. row 6) while relying on the same WRN-101 image backbone. Indeed, all multimodal methods but M3DM achieve their best performance in both tasks and setups by processing depth maps instead of point clouds. Yet, as vouched by the results achieved by PatchCore, RGB images compare favourably to depth maps. Indeed, Patchcore fed with images works much better than its depth-based counterpart and, overall, it stands out as the best baseline in our benchmark, being outperformed only by CFM (RGB + Depth) in the segmentation task dealing with the real2real setup. We ascribe this to the feature extractors used to compute depth features being pre-trained on RGB images, which may be conducive to yielding weaker features when fed with depth inputs. This points out an intriguing research question: would it be possible to develop a foundation model to process depth maps effectively?

### v2

@@ 2 Related Work
ADS Methods. Following the setup defined in the MVTec-AD benchmark, most ADS methods [22] are designed to estimate 2D anomaly maps from a single image [3, 43, 19, 29, 26, 13, 37, 27, 27, 39, 44, 34, 40, 20, 38, 23, 41, 31, 28, 16, 9, 15, 10, 2]. One of the most popular methods is PatchCore [30], which uses feature extractors pre-trained on large datasets [8, 24, 17] to gather features from nominal samples into a memory bank. At test time, input sample features are compared to this bank to detect anomalies. Another more recent yet popular approach, EfficientAD [1], uses a lightweight feature extractor and a student-teacher model to identify anomalies based on feature prediction discrepancy. The introduction of the MVTec 3D-AD benchmark has led to the development of several multimodal approaches that use images and 3D data to generate 2D anomaly maps [18, 36, 14, 32]. Among them, BTF [18] is a memory bank approach inspired by PatchCore, in which each element is the concatenation of 2D and 3D features extracted by a pre-trained image backbone and a traditional point cloud descriptor, i.e., FPFH [33], respectively. M3DM [36] improves BTF using powerful transformer-based pre-trained backbones, such as DINO-v1[8] and Point-MAE [25], to extract features from RGB images and point clouds. It also learns to fuse 2D and 3D features into multimodal features. During inference, a trained One-Class SVM learns to aggregate modality-specific scores. AST [32] follows a Teacher-Student paradigm, in which the Teacher leverages Normalizing Flows to model the multimodal features distribution, while the Student employs a feed-forward network to model such distribution. This asymmetry in the architectures exacerbates the discrepancies between nominal and anomalous features at inference time. Recently, starting from the same transformer-based backbones as in [36], CFM [14] learns to map features from one modality to the other by nominal samples and then, at inference time, detects anomalies by pinpointing inconsistencies between observed and predicted features. In this paper, we employ the above-mentioned methods to investigate the challenges of SiM3D by adapting them to work in a multiview setup.
@@ 3 Dataset Creation
Sensor. We acquired our dataset with a top-tier 3D scanner, the Atos Q from ZEISS (Fig. 3 (2.i)). It consists of a stereo pair with two 12 Mpx greyscale cameras and a light projector. We consider the left camera to be our reference. The images have a resolution of 4096×30004096$\times$ 30004096 × 3000 pixels. The point clouds can feature up to 12 million elements, with a point-to-point distance between 0.04 and 0.15 mm. The sensor is mounted on an industrial robot arm with high repeatability and precision (Fig. 3 (2.iii)).
Calibration. We calibrate the greyscale cameras with a standard chequerboard pattern, estimating the intrinsic parameters matrix, $A$, and the radial and tangential distortion coefficients, $\delta=[k_{1},k_{2},k_{3},p_{1},p_{2}]$. We also calibrate the rotation, $R_{pc}$, and translation, $T_{pc}$, from the point cloud to the left camera reference frame, with a precise dotted calibration pattern built by ZEISS (Fig. 3 (2.ii)) that provides centre coordinates with precision <1<1< 1 mm. More details are in the supplementary.
@@
Labelling. Since some anomalies, such as paintings and minor scratches, are visible only in greyscale images, while others, such as dents, are better visible in 3D, we must consider both sensing modalities to obtain precise 3D segmentation ground-truths for anomalous samples. Therefore, we develop a two-step labelling strategy, illustrated in Fig. 3 (4), that operates on both 2D images and 3D data. First, we manually annotate all the view images, $I_{i}$, where anomalies are clearly visible, obtaining dense 2D segmentation masks. We employ a different ID and colour for every defect to ease the voxel AUPRO computation (see Sec. 4). Subsequently, for each annotated image, $I_{i}$, we leverage the associated integrated mesh, $M$, the intrinsic parameters $A$, the point cloud to camera reference frame transformation, $(R_{\text{pc}},T_{\text{pc}})$, and the transformation between $v_{i}$ and $v_{1}$, $(R_{i},T_{i})$, to project the 2D annotations onto the integrated mesh. Then, since the projection may not be precise due to occlusions and slanted surfaces, each ground-truth mesh is thoroughly inspected and manually refined using a 3D visualisation and editing software, i.e., CloudCompare [12]. The labelling process was conducted by a team of 4 experts, with a final cross-check of the results. Finally, the annotated mesh is converted into a voxel grid with a voxel size of 2 mm. The choice of the size is determined by the smallest defect size discernible in the meshes. In the conversion process, a voxel is labelled as anomalous if it intersects at least one triangle labelled as such in the annotated mesh.
@@
Plastic Stool 35×35×3035$\times$ 35$\times$ 3035 × 35 × 30 330×326×317330$\times$ 326$\times$ 317330 × 326 × 317 22 1 real or 1 synth 10 10 12
Rubbish Bin 26×21×3326$\times$ 21$\times$ 3326 × 21 × 33 333×339×326333$\times$ 339$\times$ $3$ 339 × 326 42 1 real or 1 synth 20 20 12
Wicker Vase 17×17×1517$\times$ 17$\times$ 1517 × 17 × 15 329×309×290329$\times$ 309$\times$ 290329 × 309 × 290 22 1 real or 1 synth 10 10 12
Bathroom Furniture 33×33×5033$\times$ 33$\times$ 5033 × 33 × 50 371×392×363371$\times$ 392$\times$ 363371 × 392 × 363 20 1 real or 1 synth 8 10 36
Container 20×25×1020$\times$ 25$\times$ 1020 × 25 × 10 327×308×286327$\times$ 308$\times$ 286327 × 308 × 286 94 1 real or 1 synth 46 46 12
Plastic Vase 12×12×912$\times$ 12$\times$ 912 × 12 × 9 326×308×286326$\times$ 308$\times$ 286326 × 308 × 286 99 1 real or 1 synth 48 49 12
Wooden Stool 48×42×4548$\times$ 42$\times$ 4548 × 42 × 45 331×367×327331$\times$ 367$\times$ 327331 × 367 × 327 15 1 real or 1 synth 6 7 12
Sink Cabinet 44×25×5044$\times$ 25$\times$ 5044 × 25 × 50 342×349×357342$\times$ 349$\times$ 357342 × 349 × 357 19 1 real or 1 synth 9 8 36
@@ 4.1 Dataset, Tasks, and Metrics
Figure 5: Qualitatives. Selected views with defects (left) and Anomaly Volumes for several methods (right) downsampled for visualisation.
@@
Metrics. Akin to standard practice in 2D ADS [4], we evaluate detection performance by the Area Under the Receiver Operator Curve (AUROC) of the global anomaly scores predicted by the methods. As we compute a global score for each object instance, we name it instance-level AUROC (I-AUROC). Regarding segmentation performance, we evaluate methods with an adaptation of the popular Area Under the Per-Region Overlap (AUPRO) curve. First of all, we define the voxel-based Per-Region Overlap (PRO) metric as $\text{PRO}_{t}=\frac{1}{N}\sum_{n=1}^{N}\frac{|\overline{\Psi}_{t}\cap V_{n}|}{|V_{n}|}$ , where $\overline{\Psi}_{t}$ is an Anomaly Volume binarised using a threshold, $t$, $N$ is the number of blobs, i.e., defects, in the ground-truth, and $V_{n}$ are the voxels belonging to that blob. Since during labelling, we annotated each defect in our dataset with a different ID (see Sec. 3), the $N$ defect blobs in a ground-truth are easily identifiable. This metric is computed across multiple thresholds, $t$, estimated at different False Positive Rates (FPRs). We calculate the FPR using only non-empty voxels. We sample FPRs from 0 to the maximum value uniformly, we estimate the corresponding threshold $t$, and we binarise the score grid accordingly. In this way, we construct a curve that plots the PRO values against the corresponding FPRs. The segmentation performance is obtained by integrating such a curve up to a specific False Positive Rate. As proposed in [14], we use 1% as integration bound for the PRO curve to set a challenging bar representative of the requirements of real industrial applications. Finally, we normalise the resulting area to the interval [0,1][0,1][ 0 , 1 ]. We dub this metric voxel-level AUPRO (V-AUPRO).
@@ 4.2 Baselines
Extending singleview methods to Multiview 3D ADS. Current ADS methods predict a 2D anomaly map from a singleview input, either unimodal, e.g., images, or multimodal, e.g., images and point clouds. To obtain a set of baselines for our novel multiview 3D ADS task, we extend popular and effective singleview methods by means of the 3D aggregation strategy illustrated in Fig. 4. Given the considered method, we process the data corresponding to an input view, $v_{i}$, so as to obtain a 2D Anomaly Map. Similarly to Sec. 3, given the mesh, $M$, intrinsic parameters $A$, point cloud to camera reference frame transformation, $(R_{\text{pc}},T_{\text{pc}})$ and view pose $v_{0}$, $(R_{i},T_{i})$, we can project each pixel of the 2D Anomaly Map to its corresponding 3D position. We discretise the 3D position, finding the corresponding voxel index. We associate that voxel with the corresponding anomaly score. We repeat the process for each view, keeping track of the scores that project onto the same voxels. Eventually, we compute the maximum among all scores projected into each voxel, obtaining the final Anomaly Volume. To extract the global score required by the I-AUROC, we compute the maximum across all the scores within the Anomaly Volume.
@@
Selected Methods. We extend some of the most popular and effective unimodal and multimodal singleview ADS methods, starting from their official code, and evaluate them on SiM3D. Regarding unimodal methods, we select PatchCore [30], with either WideResNet101 [42] or DINO-v2 [24] as backbone, to extract features from images of all views, $\{I_{i}\}_{i=1}^{N}$ of the single-instance training set to create memory banks. Moreover, we use EfficientAD [1] using all views as their training set. As for multimodal ADS methods, we use BTF [18] employing WideResNet101 [42] as the image encoder and FPFH [33] as the algorithm to extract point cloud features. We create memory banks using features extracted from images, $\{I_{i}\}_{i=1}^{N}$, and point clouds, $\{P_{i}\}_{i=1}^{N}$, of all views of the single-instance training set. We use the same strategy with M3DM [36]. As 2D feature extractor, we exploit DINO-v2. As point cloud feature extractor, instead of using a transformer backbone, i.e., Point-MAE [17], we utilise FPFH to extract 3D features, similarly to BTF. Indeed, we would need a severe downsample (from $\sim$5 – 7M to $\sim$8k points) to work with Point-MAE, and we verified that at such low resolution, most of the defects no longer show up in the point clouds. Finally, we select CFM [14], using all views and clouds as a single training set. Based on the same consideration as for M3DM, we use DINO-v2 and FPFH to produce 2D and 3D features. Due to computational constraints, all methods process images padded and downsampled to $\times 1540$ 1540. Moreover, for compatibility with the employed 2D feature extractors, which accept RGB images, we create 3-channel inputs by stacking our greyscale images thrice.
@@
3D Data pre-processing. To obtain less noisy data and improve the results of the above-mentioned baselines, we remove the background from 3D data. We estimate a 3D plane using a RANSAC-based plane segmentation algorithm, which, iteratively, randomly selects a small subset of points (e.g., 10) to fit a candidate plane model. For the candidate planes, we compute the consensus set with threshold $\tau$ on the point-to-plane distance. After 1000 iterations, we return the best-fitting plane. To account for spurious background artefacts, we shift the plane by an offset $\alpha$ and finally remove all points below the shifted plane. Given the different sizes of the objects, for each object type, we select different optimal thresholds $\tau$ and offsets $\alpha$, as detailed in the supplementary material. We apply this background removal technique to the integrated mesh, $M$, due to the larger number of background points, which facilitates the RANSAC algorithm. Then, from the background-filtered integrated mesh, we leverage ray-tracing techniques and known camera parameters to render, for each view, depth maps, $\{D_{i}\}_{i=1}^{N}$, pixel-aligned to the corresponding images, $\{I_{i}\}_{i=1}^{N}$ at the sensor resolution (4096×30004096$\times$ 30004096 × 3000). Finally, these depth maps are padded and downsampled to $\times 1540$ 1540, and the depth is lifted in 3D via the camera parameters to obtain dense and clean point clouds. The depth maps and point clouds obtained by rendering the pre-processed meshes serve as the input data for all the experiments presented in the subsequent section and will be available in the SiM3D dataset.
@@ 5 Benchmark Results
Main Results. In  Tab. 3 and  Tab. 4, we report the results obtained on the SiM3D benchmark by extending the considered methods as Multiview 3D ADS baselines. By looking at the mean detection performance (Tab. 3), we notice how methods that natively process only RGB images outperform natively multimodal ones in both the real2real and synth2real setups. With the notable exception of AST, this also stands for mean segmentation performance (Tab. 4). We posit that this is due to the main multimodal methods having been designed and tuned to process point clouds of the resolution featured by existing multimodal benchmarks, which is much lower than that of SiM3D, i.e. $\sim$ 8K vs $\sim$ 2.3M points (after downsampling). Indeed, AST is the only one designed to work natively with depth maps. Hence, our experiments suggest that current multimodal ADS methods do not scale to significantly higher resolutions. In turn, we could not even deploy the native 3D feature extractors of M3DM and CFM, the methods currently excelling on [6, 7]. Furthermore, we notice how methods relying on memory banks, such as PatchCore and M3DM, tend to attain slightly better mean detection and segmentation performance than their counterparts within the same modality that require substantial training, i.e., EfficientAD and CFM. We argue that this may be ascribed to the single-instance setup, which renders the latter category of approaches harder to optimise due to the limited number of training images, which are as many as the views. The methods most effectively extended to our multiview 3D task are Patchcore, which, equipped with either WRN-101 or DINO-v2, can achieve the best detection in both setups, and AST, achieving the best segmentation performance in both setups. Indeed, Patchcore yields I-AUROCs of 0.754 (WRN-101) and 0.540 (DINO-v2), while AST yields a V-AUPRO@1% of 0.699 and 0.695, in the real2real and synth2real setups, respectively. It is also worth pointing out how, due to the domain shift, we can notice a gap between the real2real and synth2real setups, especially in detection (0.754 vs. 0.540, I-AUROC) compared to segmentation performance (0.699 vs. 0.695, V-AUPRO@1%). The qualitative results, shown in Fig. 5, visually support the above observations. Overall, the relatively low, definitely far from saturated, performance figures provided by baselines designed by straightforward extension of state-of-the-art methods delivering impressive results in singleview 2D benchmarks hint at the novel and challenging nature of the task set forth by SiM3D. Thus, our experimental findings by baseline approaches reveal the need for more advanced and specific methods and paradigms, both in the real2real and, even more so, the synth2real setups.
@@
Alternative 3D Representations. Given the lack of feature extractors amenable to processing high-resolution point clouds, we explore the use of depth maps to capture 3D cues. The 2D grid structure of depth maps enables the deployment of popular backbones pre-trained on RGB data that support high-resolution inputs, such as DINO-v2. Thus, we create and evaluate additional baselines using depth maps and report the results in  Tab. 5. Similarly to greyscale images, we tile depth maps along the channel axis to ensure compatibility with the employed backbones. The results show that the use of depth maps rather than point clouds in multimodal methods tends to improve detection and segmentation performance by sizable margins, e.g., in the real2real setup, BTF yields 0.707 I-AUROC and 0.609 V-AUPRO when processing depths compared to 0.478 I-AUROC and 0.444 V-AUPRO when fed with point clouds (row 6 vs. row 7) while relying on the same WRN-101 image backbone. Indeed, all multimodal methods but M3DM achieve their best performance in both tasks and setups by processing depth maps instead of point clouds. Yet, as vouched by the results achieved by PatchCore, RGB images compare favourably to depth maps in terms of detection performance. We ascribe this to the feature extractors used to compute depth features being pre-trained on RGB images, which may be conducive to yielding weaker features when fed with depth inputs. This points out an intriguing research question: would it be possible to develop a foundation model to process depth maps effectively?